Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle?
Answer by checkley77(12569)

600/X-600/(X+20)=1
(600[X+20]-600*X)/X(X+20)=1
(600X+12000-600X)/(X^2+20X)=1 NOW CROSS MULTIPLY
X^2+20X=12,000
X^2+20X-12,000=0
(X+120)(X-100)=0
X-100=0
X=100 SPEED OF THE TRIP.
PROOF:
600/100-600/120=1
6-5=1
1=1